# WaveNet’s Precision in EEG Classification

### v1

@@ Abstract
This study introduces a WaveNet-based deep learning model designed to automate the classification of EEG signals into physiological, pathological, artifact, and noise categories. Traditional methods for EEG signal classification, which rely on expert visual review, are becoming increasingly impractical due to the growing complexity and volume of EEG recordings. Leveraging a publicly available annotated dataset from Mayo Clinic and St. Anne’s University Hospital, the WaveNet model was trained, validated, and tested on 209,232 samples with a 70/20/10 % split. The model achieved a classification accuracy exceeding previous CNN and LSTM-based approaches, and was benchmarked against a Temporal Convolutional Network (TCN) baseline. Notably, the model distinguishes noise and artifacts with high precision, although it reveals a modest but explainable degree of misclassification between physiological and pathological signals, reflecting inherent clinical overlap. WaveNet’s architecture, originally developed for raw audio synthesis, is well-suited for EEG data due to its use of dilated causal convolutions and residual connections, enabling it to capture both fine-grained and long-range temporal dependencies. The research also details the preprocessing pipeline, including dynamic dataset partitioning and normalization steps that support model generalization.
@@ I Introduction
The classification of EEG data into physiological, pathological, and noise categories is essential for accurate neurophysiological analysis and clinical diagnostics. Traditional methods rely heavily on expert visual review, which is not scalable given the increasing volume and complexity of EEG recordings. Trained EEG technologists are crucial in identifying abnormalities and enhancing patient outcomes. Manual EEG review by neurophysiologists and technologist is prone to errors, non-scalability, subjectivity, and inconsistencies due to varying expertise, fatigue, and cognitive biases [1]. In contrast, automated models [2, 3] are scalable, mitigate oversight, and use deep learning (DL) to discern subtle EEG patterns, improving reliability for precise diagnoses and treatments.
@@
A Temporal Convolutional Network (TCN) is a type of DL neural network designed for processing sequential data. WaveNet is a specific type of TCN that excels at handling the intricate temporal characteristics of audio waveforms due to its architecture, which includes dilated causal convolutions and residual connections [4]. In WaveNet, causal convolutions ensure that the output at a time step depends on the current and previous time steps, mimicking the autoregressive nature of time series data like EEG signals. This is crucial as it prevents information leakage from future time steps into the current prediction.
@@
EEG data are continuous signals reflecting the brain’s electrical activity, exhibiting complex temporal dependencies and dynamic variations [5]. The architecture is particularly well-suited for this task, as it can capture both fine-grained and long-range temporal dependencies in the data. This capability makes WaveNet a strong option for EEG data classification:
@@
  • Temporal Dynamics: WaveNet’s ability to model long-range dependencies and temporal patterns aligns well with capturing EEG signals’ temporal dynamics. EEG signals fit this description perfectly, as they exhibit complex temporal dependencies and variations (see Figure 1 and 2 for examples).
@@
  • Hierarchical Features: WaveNet’s hierarchical nature allows it to learn multi-scale features, crucial for identifying patterns in EEG data at different temporal resolutions. High-resolution features are captured by lower layers, while higher layers capture long-term, low-resolution patterns. This hierarchical feature extraction is crucial for modeling complex temporal dynamics in sequential data, which can allow the model to differentiate between physiological and pathological EEG data (Figures 1 and 2) and noise or artifacts.
@@
  • High Resolution: WaveNet can capture detailed variations in EEG signals, essential for accurate classification, like the differences in the peaking behaviour of the EEG signals of pathological and healthy patients (Fig. 1 and Fig. 2).
@@
Therefore, we propose using a WaveNet model to automate the classification of EEG data. To provide context, we also benchmark its performance against a TCN baseline. WaveNet’s architecture makes it well-suited for capturing the complex temporal patterns in EEG signals. Automating EEG data classification with WaveNet will not only enhance efficiency and accuracy but also ensure more consistent and reliable diagnostics, ultimately improving patient outcomes and advancing the field of neurophysiology.
@@
Figure 2: Example of abnormal brain activity observed in epilepsy-prone neurons is identified by a sharp wave transient. Indicated by the red arrow one can see this abnormal transient, a high frequency oscillation (HFO) can be observed riding atop the peak of the spikes.
@@ II Related works
Accurate detection of epileptiform spikes is essential for the diagnosis and clinical management of epilepsy. Earlier methods relied heavily on manual annotation or rule-based algorithms, which are time consuming and subject to inter-rater variability, while recent deep learning approaches show promise in automating EEG spike classification by learning features directly from raw or minimally processed signals [6].
Initial work applied convolutional neural networks (CNNs) to this task. Johansen et al. [7] reported a CNN-based classifier with an area under the curve (AUC) of 95 %, effectively identifying epileptiform spikes with high sensitivity and low false positive rates. Nejedly et al. [3] extended this approach by combining graphoelement-derived features with both CNN and long short-term memory (LSTM) architectures, achieving F1 scores of 0.81 and 0.80, respectively. Their specialized patient tuned CNN model for pathological activity classification reached an average F1 score of 0.96. They also explored artifact rejection in intracerebral EEG using CNN-based classifiers [2], improving reliability by filtering out physiological noise from the input signals.
Subsequent research incorporated attention mechanisms into recurrent models to better capture temporal dependencies. Fukumori, Yoshida, and Tanaka [8] proposed a self-attentive recurrent neural network that dynamically emphasized relevant time steps, achieving an average accuracy of 90 % in epileptic spike detection. Similarly, Rukhsar and Tiwari [9] introduced an Attention Recurrent Neural Network (ARNN) that applied self-attention across multi-channel EEG data. Evaluated on datasets like CHB-MIT and canine EEG from the Mayo Clinic and North Carolina State University, the model achieved an average F1 score of 0.94 and outperformed LSTM and Vision Transformer baselines. However, ARNN remains a specialized model trained and evaluated within specific datasets, and its performance on cross-patient or cross-institution data remains an open question.
@@ III Dataset
The dataset for this article comprises annotated intracranial EEG recordings from Mayo Clinic (MN, USA) and St. Anne’s University Hospital (Brno, Czech Republic). The dataset includes EEG data labelled into three categories: physiological activity, pathological/epileptic activity, and artifactual signals. Each data sample of 3 seconds is stored in a .mat file with a sampling frequency of 5kHz and consists of a vector with dimensions $(1,15000)$ containing a total of 209,232 samples. Comprehensive instructions for dataset usage are provided in the associated scientific papers [2][3].
@@ IV-A Preprocessing steps
The preprocessing steps involved loading EEG data from .mat files stored in directories, and organizing them into HDF5 format. This format optimizes storage and access speed while handling large datasets efficiently. After this preprocessing step, the data is dynamically split into distinct training, validation, and test sets. This splitting is achieved through random sampling of the data keys, ensuring separation and preventing data leakage between the sets during model evaluation and training. The entire pipeline is designed with mechanisms to maintain the integrity and separation of data throughout each stage of processing. Then, the input layer normalizes the input using Z-score normalization, adding $\varepsilon=1\times 10^{-8}$ to prevent division by zero.
@@ IV-B Model structure
The WaveNet model is built using dilated causal convolutions to capture long-range dependencies in EEG signals. It employs seven sequential residual blocks with increasing dilation rates of $[1,2,4,8,16,32,64]$, allowing the receptive field to grow exponentially while maintaining the temporal ordering for time series modeling.
@@
These skip connections are summed and passed through two additional Conv1D layers. The first with 32 filters, and the second with 4 filters (equal to the number of output classes). All before being globally averaged across the time dimension via a GlobalAveragePooling1D layer. A final softmax activation layer produces the class probability distribution for each EEG segment.
To combat overfitting, an adaptive dropout mechanism is employed. This starts at an initial rate of 0.20 and dynamically adjusts throughout training based on a composite score combining validation accuracy, AUC, loss, and the relative gap between training and validation performance (for more detail see appendix A). The final observed dropout rate converged to 0.23 by the end of training.
Optimization is handled by the Adam [10] optimizer with an initial learning rate of 0.001, and L2 regularization with $\lambda=0.0001$ is applied to constrain model complexity and enhance generalization. To address class imbalance, particularly the dominance of physiological and pathological EEG patterns, the training pipeline applies dynamic class weighting called focal loss[11]. When using focal loss, class weights inversely proportional to class frequencies are passed as alpha parameters, focusing learning on underrepresented classes and down-weighting well-classified samples.
@@ IV-C Baseline: Temporal Convolutional Network (TCN)
The Temporal Convolutional Network (TCN) was introduced as an additional baseline to provide a convolutional, non-autoregressive counterpart to WaveNet. While both architectures employ causal, dilated convolutions and residual connections, the TCN is tailored for classification rather than autoregressive synthesis. Its design kept receptive field and capacity comparable to WaveNet.
Architecture:
@@ IV-D Training summary
The wavenet model was trained on the Kaggle platform using a Tesla P100 GPU with optimized hardware acceleration. A custom focal loss [11] function addressed class imbalance, while adaptive dropout and adam were used to improve generalization. Key hyperparameters, including batch size, learning rate, and dilation settings, were carefully tuned for optimal performance. For a comprehensive overview of the training configuration and hardware specifics, see appendix A.
The TCN model training was conducted locally on a MSI Katana laptop using the NVIDIA GeForce RTX 3050 Laptop GPU with mixed-precision acceleration. A custom focal loss function [11] was used to mitigate class imbalance, while Adam optimization and a fixed dropout of 0.005 within residual blocks supported generalization. Key hyperparameters, including dilation rates, batch size (16), learning rate (1e-3), and number of filters (8), were tuned empirically for optimal performance.
@@ V Results
The WaveNet model was trained on a curated dataset from St. Anne’s University Hospital and the Mayo clinic consisting of 292,232 EEG samples, split into 70% training, 20% validation, and 10% testing subsets. The model achieved an average F1 score of 0.94, outperforming prior methods such as the general CNN approach [2] and the graphoelements Convolutional LSTM model [3] which obtained an average F1 scores of 0.80 and 0.81 respectively (see Table III). While underperforming on the specialized model which reached a F1 of 0.96.
@@
Figure 3: Accuracy figure, plotted every 150 batches each batch contains 32 samples for the training set. Yellow for the validation set plotted every epoch.
Figure 4: Validation and training loss figure, plotted every 150 batches each batch contains 64 samples for the training set. Yellow for the validation set plotted every epoch.
@@
Overall, these results highlight the WaveNet model’s balance of sensitivity and specificity across classes, achieving precision and recall values exceeding 0.93 for most categories (Table III). This positions the model as a powerful tool for clinical EEG analysis, with clear advantages in suppressing irrelevant variance while preserving meaningful neurophysiological signals.
VI Conclusion
This study demonstrates that the WaveNet architectures, originally developed for audio synthesis, can be effectively repurposed for EEG classification. Its capacity to capture both fine-grained waveform subtleties and long-range temporal dependencies makes it well-suited for neurophysiological data. Leveraging a merged dataset from St. Anne’s University Hospital (Brno, Czech Republic) and the Mayo Clinic (Rochester, Minnesota, USA), the model achieved high validation and test accuracies, despite significant class imbalance and inter-subject variability.
Crucially, the use of dilated convolutions enabled the network to represent scale-variant features and rare event types, key characteristics of EEG and other biological signals. This temporal expressiveness, combined with adaptive regularization strategies such as dynamic dropout and focal loss with class reweighing, enhanced generalizability. The model’s strong cross-clinic performance suggests viability for real-world applications including seizure detection, artifact rejection, and closed-loop brain-computer interfaces.
WaveNet outperformed generalized CNN models across most metrics, while the specialized CNN achieved higher performance on physiological EEG signals (F1 0.96 vs. 0.98; Table III). The TCN baseline trained on the same preprocessing and dataset splits achieved performance comparable to WaveNet on several classes, demonstrating that a non-autoregressive convolutional architecture can capture EEG temporal patterns effectively. The specialized CNN reported by Nejedly et al. [2] highlights the trade off of highly tuned models: they maximize performance on particular signal types but require patient-specific tuning. WaveNet provides strong, consistent performance across classes, including improved pathological classification relative to the generalized CNN (F1 0.81 vs. 0.64). Which might reflect its deeper temporal receptive field and residual connections that capture long-range or irregular dynamics in EEG signals.
Notably, the WaveNet model was trained without patient specific tuning, yet achieved strong performance across clinically diverse EEG datasets. This outcome shows its capacity to generalize under real world conditions marked by variability and imbalance. Such inherent flexibility points toward a broader application in synthesize long range temporal dependencies and multi channel interactions in high density neural recordings. In particular, it opens the door to applying similar architectures to Neuropixels data within spike sorting pipelines like Kilosort, where waveform diversity and spatio-temporal structure require adaptable and scalable models. Building on these principles, future work will extend this approach to neuronal spike classification in large-scale electrophysiological recordings.
@@ VI Conclusion
Additionally, while the WaveNet model demonstrated cross-hospital generalization on diverse human EEG datasets, other architectures, such as the recently proposed Attention Recurrent Neural Network (ARNN) by Rukhsar and Tiwari [9], have reported impressive accuracy on long-duration, multi-channel seizure datasets, including CHB-MIT and canine EEG from the Mayo Clinic and North Carolina State University. While ARNN achieves strong accuracy, particularly in per-subject evaluations, it utilized 30 epochs for training, whereas our WaveNet model’s training had a hard stop at 10 epochs due to limited computing access. Future research should directly compare WaveNet and ARNN models on shared, heterogeneous EEG datasets, under similar training constraints, to evaluate their respective trade-offs in generalization, computational efficiency, and interpretability.
@@
Both models leverage temporal pattern recognition for their high computational efficiency and accuracy. The ARNN model uses an LSTM-style recurrent gate to process sequences, while our WaveNet model employs dilated causal convolutions for temporal pattern recognition, followed by two convolutional ending layers. Given that traditional LSTM models, as baselines in other studies, have shown comparatively lower performance than more advanced temporal models on certain EEG datasets, an intriguing future direction is to explore hybrid architectures. Specifically, combining WaveNet’s unique capacity for capturing long-range temporal dependencies via dilated convolutions with the sequence modeling capabilities of recurrent networks could potentially enhance the efficiency and accuracy of models like ARNN.

### v2

# WaveNet’s Precision in iEEG Classification

@@ Abstract
This study introduces a WaveNet-based deep learning model designed to automate the classification of Intracranial Electroencephalography (iEEG) signals into physiological activity, pathological/epileptic activity, power-line noise, and other non-cerebral artifacts categories. Traditional methods for iEEG signal classification, which rely on expert visual review, are becoming increasingly impractical due to the growing complexity and volume of iEEG recordings. Leveraging a publicly available annotated dataset from Mayo Clinic and St. Anne’s University Hospital, the WaveNet model was trained, validated, and tested on 209,231 samples with a 70/20/10 % split. The model achieved a classification accuracy exceeding previous non specialized CNN and LSTM-based approaches, and was benchmarked against a Temporal Convolutional Network (TCN) baseline. Notably, the model achieves high discrimination of noise and artifact classes (precision = 0.98 and $\approx 1$, respectively). Although, classification between physiological and pathological signals exhibits a modest but clinically interpretable overlap, with F1-scores of 0.96 and 0.90 and 175 and 272 cross-class false positives, respectively. Reflecting inherent clinical overlap. WaveNet’s architecture, originally developed for raw audio synthesis, is well-suited for iEEG data due to its use of dilated causal convolutions and residual connections, enabling it to capture both fine-grained and long-range temporal dependencies. The research also details the preprocessing pipeline, including dynamic dataset partitioning, the use of focal loss to combat class imbalances and normalization steps that support model high performance. While results demonstrate strong in-distribution performance, generalizability across datasets and clinical settings has yet to be established.
@@ I Introduction
The classification of Intracranial Electroencephalography (iEEG) data into physiological, pathological, artifact, and noise categories is essential for accurate neurophysiological analysis and clinical diagnostics. Traditional methods rely heavily on expert visual review, which is not scalable given the increasing volume and complexity of iEEG recordings. Trained iEEG technologists are crucial in identifying abnormalities and enhancing patient outcomes. Manual EEG review by neurophysiologists and technologist is prone to errors, non-scalability, subjectivity, and inconsistencies due to varying expertise, fatigue, and cognitive biases [6]. In contrast, automated models [9, 10] are scalable, mitigate oversight, and use deep learning (DL) to discern subtle iEEG patterns, improving reliability for precise diagnoses and treatments.
@@
A Temporal Convolutional Network (TCN) is a type of DL neural network designed for processing sequential data. WaveNet is a specific type of TCN that excels at handling the intricate temporal characteristics of audio waveforms due to its architecture, which includes dilated causal convolutions and residual connections [12]. In WaveNet, causal convolutions ensure that the output at a time step depends on the current and previous time steps, mimicking the autoregressive nature of time series data like iEEG signals. This is crucial as it prevents information leakage from future time steps into the current prediction.
@@
iEEG data are continuous signals reflecting the brain’s electrical activity, exhibiting complex temporal dependencies and dynamic variations [3]. The architecture is particularly well-suited for this task, as it can capture both fine-grained and long-range temporal dependencies in the data. This capability makes WaveNet a strong option for iEEG data classification:
@@
  • Temporal Dynamics: WaveNet’s ability to model long-range dependencies and temporal patterns aligns well with capturing iEEG signals’ temporal dynamics. iEEG signals fit this description perfectly, as they exhibit complex temporal dependencies and variations (see Figure 1 and 2 for examples).
@@
  • Hierarchical Features: WaveNet’s hierarchical nature allows it to learn multi-scale features, crucial for identifying patterns in iEEG data at different temporal resolutions. High-resolution features are captured by lower layers, while higher layers capture long-term, low-resolution patterns. This hierarchical feature extraction is crucial for modeling complex temporal dynamics in sequential data, which can allow the model to differentiate between physiological and pathological iEEG data (Figures 1 and 2) and noise or artifacts.
@@
  • High Resolution: WaveNet can capture detailed variations in iEEG signals, essential for accurate classification, like the differences in the peaking behaviour of the iEEG signals of pathological and healthy patients (Fig. 1 and Fig. 2).
@@
Therefore, we propose using a WaveNet model to automate the classification of iEEG data. To provide context, we also benchmark its performance against a TCN baseline. WaveNet’s architecture makes it well-suited for capturing the complex temporal patterns in iEEG signals. Automating iEEG data classification with WaveNet will not only enhance efficiency and accuracy but also ensure more consistent and reliable diagnostics, ultimately improving patient outcomes and advancing the field of neurophysiology.
  • Temporal Dynamics: Long-range dependencies and temporal patterns in iEEG.
  • Hierarchical Features: Multi-scale features, distinguishing physiological, pathological, and artifact signals.
  • High Resolution: Detailed waveform variations crucial for precise classification.
@@
Figure 2: Example of abnormal brain activity observed in epilepsy-prone neurons identified by a sharp wave transient. High frequency oscillation (HFO) observed atop the spike peak.
@@ II Related works
Accurate detection of epileptiform spikes is essential for the diagnosis and clinical management of epilepsy. Earlier methods relied heavily on manual annotation or rule-based algorithms, which are time consuming and subject to inter-rater variability, while recent deep learning approaches show promise in automating (i)EEG spike classification by learning features directly from raw or minimally processed signals [1].
Initial work applied convolutional neural networks (CNNs) to this task. Johansen et al. [4] reported a CNN-based classifier with an area under the curve (AUC) of 95 %, effectively identifying epileptiform spikes with high sensitivity and low false positive rates. Nejedly et al. [10] extended this approach by combining graphoelement-derived features with both CNN and long short-term memory (LSTM) architectures, achieving F1 scores of 0.81 and 0.80, respectively. Their specialized patient tuned CNN model for pathological activity classification reached an average F1 score of 0.96. They also explored artifact rejection in iEEG using CNN-based classifiers [9], improving reliability by filtering out physiological noise from the input signals.
Subsequent research incorporated attention mechanisms into recurrent models to better capture temporal dependencies. Fukumori, Yoshida, and Tanaka [2] proposed a self-attentive recurrent neural network that dynamically emphasized relevant time steps, achieving an average accuracy of 90 % in epileptic spike detection. Similarly, Rukhsar and Tiwari [11] introduced an Attention Recurrent Neural Network (ARNN) that applied self-attention across multi-channel iEEG data. Evaluated on datasets like CHB-MIT and canine iEEG from the Mayo Clinic and North Carolina State University, the model achieved an average F1 score of 0.94 and outperformed LSTM and Vision Transformer baselines. However, ARNN remains a specialized model trained and evaluated within specific datasets, and its performance on cross-patient or cross-institution data remains an open question.
@@ III Dataset
The dataset for this article comprises annotated iEEG recordings from Mayo Clinic (MN, USA) and St. Anne’s University Hospital (Brno, Czech Republic). The dataset includes iEEG data labelled into four categories: physiological activity, pathological/epileptic activity, power-line noise, and other non-cerebral artifacts. Each data sample of 3 seconds is stored in a .mat file with a sampling frequency of 5kHz and consists of a vector with dimensions $(1,15000)$ containing a total of 209,231 samples. Comprehensive instructions for dataset usage are provided in the associated scientific papers [9][10].
@@ IV-A Preprocessing steps
The preprocessing steps involved loading iEEG data from .mat files stored in directories, and organizing them into HDF5 format. This format optimizes storage and access speed while handling large datasets efficiently. After this preprocessing step, the data is dynamically split into distinct training, validation, and test sets. This splitting is achieved through random sampling individual samples, ensuring separation and preventing data leakage between the sets during model evaluation and training. The entire pipeline is designed with mechanisms to maintain the integrity and separation of data throughout each stage of processing. Then, the input layer normalizes the input using Z-score normalization, adding $\varepsilon=1\times 10^{-8}$ to prevent division by zero.
@@ IV-B Model structure
The WaveNet model is built using dilated causal convolutions to capture long-range dependencies in iEEG signals. It employs seven sequential residual blocks with increasing dilation rates of $[1,2,4,8,16,32,64]$, allowing the receptive field to grow exponentially while maintaining the temporal ordering for time series modeling.
@@
These skip connections are summed and passed through two additional Conv1D layers. The first with 32 filters, and the second with 4 filters (equal to the number of output classes). All before being globally averaged across the time dimension via a GlobalAveragePooling1D layer. A final softmax activation layer produces the class probability distribution for each iEEG segment.
To combat overfitting, an novel adaptive dropout mechanism is employed. This starts at an initial rate of 0.20 and dynamically adjusts throughout training based on a composite score combining validation accuracy, AUC, loss, and the relative gap between training and validation performance (for more detail see appendix A). The final observed dropout rate converged to 0.23 by the end of training.
Optimization is handled by the Adam [5] optimizer with an initial learning rate of 0.001, and L2 regularization with $\lambda=0.0001$ is applied to constrain model complexity and enhance generalization. To address class imbalance, particularly the dominance of physiological and pathological iEEG patterns, the training pipeline applies dynamic class weighting called focal loss[7]. When using focal loss, class weights inversely proportional to class frequencies are passed as alpha parameters, focusing learning on underrepresented classes and down-weighting well-classified samples.
@@ IV-C Baseline: Temporal Convolutional Network (TCN)
The Temporal Convolutional Network (TCN) was introduced as an additional baseline to provide a convolutional, non-autoregressive counterpart to WaveNet. While both architectures employ causal, dilated convolutions and residual connections, the TCN is tailored for classification rather than autoregressive synthesis. Its design kept receptive field and capacity comparable to WaveNet. Architecture:
@@ IV-D Training summary
The WaveNet and TCN models were trained on different hardware platforms but shared a common optimization strategy. Both models employed a custom focal loss function [7] to address class imbalance and used Adam optimization with dropout to mitigate overfitting.
WaveNet was trained on the Kaggle platform using a Tesla P100 GPU with optimized hardware acceleration and adaptive dropout, with batch size, learning rate, 36,710 trainable parameters and 2 non trainable parameters (see Appendix A for full details).
TCN training was conducted locally on an MSI Katana laptop equipped with an NVIDIA GeForce RTX 3050 Laptop GPU using mixed-precision acceleration; fixed dropout (0.005) was applied within residual blocks, and dilation rates, batch size (16), learning rate (1e-3), and number of filters (8) were empirically optimized. It contained 4,852 parameters in total, of which 4,402 are trainable.
@@ V Results
The WaveNet model was trained on a curated dataset from St. Anne’s University Hospital and the Mayo clinic consisting of 209,231 EEG samples, split into 70% training, 20% validation, and 10% testing subsets. The model achieved an average F1 score of 0.94, outperforming prior methods such as the general CNN approach [9] and the graphoelements Convolutional LSTM model [10] which obtained an average F1 scores of 0.80 and 0.81 respectively (see Table III). While underperforming on the specialized model which reached a F1 of 0.96.
@@
Figure 3: Wavenet accuracy figure, plotted every 150 batches each batch contains 32 samples for the training set. Yellow for the validation set plotted every epoch Minor spikes at epoch boundaries reflect the “epochal sawtooth phenomenon” [8], caused by data reshuffling and optimizer dynamics.
Figure 4: Wavenet validation and training loss figure, plotted every 150 batches each batch contains 32 samples for the training set. Yellow for the validation set plotted every epoch. Transient spikes at the start of epochs correspond to the same Epochal Sawtooth phenomenon as explained by liu et al(2025).
@@
Overall, these results highlight the WaveNet model’s balance of sensitivity and specificity across classes, achieving precision and recall values exceeding 0.93 for most categories (Table III). This positions the model as a powerful tool for clinical iEEG analysis, with clear advantages in suppressing irrelevant variance while preserving meaningful neurophysiological signals.
VI Discussion
Notably, the WaveNet model was trained without patient specific tuning, yet achieved strong performance across clinically diverse iEEG datasets. This outcome shows its capacity to generalize under real world conditions marked by variability and imbalance. Such inherent flexibility points toward a broader application in synthesize long range temporal dependencies and multi channel interactions in high density neural recordings. In particular, it opens the door to applying similar architectures to Neuropixels data within spike sorting pipelines like Kilosort, where waveform diversity and spatio-temporal structure require adaptable and scalable models. Building on these principles, future work will extend this approach to neuronal spike classification in large-scale iEEG recordings.
@@ VI Discussion
Additionally, while the WaveNet model demonstrated high performance on diverse human iEEG datasets, other architectures, such as the recently proposed Attention Recurrent Neural Network (ARNN) by Rukhsar and Tiwari [11], have reported impressive accuracy on long-duration, multi-channel seizure datasets, including CHB-MIT and canine iEEG from the Mayo Clinic and North Carolina State University. While ARNN achieves strong accuracy, particularly in per-subject evaluations, it utilized 30 epochs for training, whereas our WaveNet model’s training had a hard stop at 10 epochs due to limited computing access. Future research should directly compare WaveNet and ARNN models on shared, heterogeneous iEEG datasets, under similar training constraints, to evaluate their respective trade-offs in generalization, computational efficiency, and interpretability.
@@
Both models leverage temporal pattern recognition for their high computational efficiency and accuracy. The ARNN model uses an LSTM-style recurrent gate to process sequences, while our WaveNet model employs dilated causal convolutions for temporal pattern recognition, followed by two convolutional ending layers. Given that traditional LSTM models, as baselines in other studies, have shown comparatively lower performance than more advanced temporal models on certain iEEG datasets, an intriguing future direction is to explore hybrid architectures. Specifically, combining WaveNet’s unique capacity for capturing long-range temporal dependencies via dilated convolutions with the sequence modeling capabilities of recurrent networks could potentially enhance the efficiency and accuracy of models like ARNN.
VII Limitations
Despite the high classification performance of the WaveNet model, several limitations must be acknowledged. First, the model’s performance on the distinction between physiological and pathological iEEG signals, while high, is not absolute. The observed cross-class false positives (175 for physiological and 272 for pathological) suggest that the model occasionally struggles with the inherent morphological similarity between some healthy sharp transients and epileptiform activity. This reflects a broader challenge in clinical neurophysiology where definitive ground truth can sometimes be subjective.
Second, the current study utilized a fixed 3-second windowing approach. While effective for capturing the transient nature of spikes and noise, it may not fully capture the slower, evolutional temporal dynamics of longer ictal events or state-dependent changes in the background EEG. Third, due to computational resource constraints, the training was limited to 10 epochs. While the convergence curves (Fig. 3 and 4) suggest stability, a more exhaustive hyperparameter search and extended training could potentially yield further marginal gains in precision.
Finally, while the dataset includes multi-site data from the Mayo Clinic and St. Anne’s University Hospital, the model’s performance on completely ”out-of-distribution” data ,such as recordings from different amplifier hardware or significantly different sampling rates, remains untested. Future work should focus on cross-dataset validation to ensure the WaveNet’s dilated convolutions have learned generalizable biomarkers rather than site-specific signatures.
VIII Conclusion
This study demonstrates that the WaveNet architectures, originally developed for audio synthesis, can be effectively repurposed for classification of physiological, pathological, artifact, and noise in iEEG . Its capacity to capture both fine-grained waveform subtleties and long-range temporal dependencies makes it well-suited for neurophysiological data. Leveraging a merged dataset from St. Anne’s University Hospital (Brno, Czech Republic) and the Mayo Clinic (Rochester, Minnesota, USA), the model achieved high validation and test accuracies, despite significant class imbalance and inter-subject variability.
WaveNet outperformed generalized CNN models across most metrics, while the specialized CNN achieved higher performance on physiological iEEG signals (F1 0.96 vs. 0.98; Table III). The TCN baseline trained on the same preprocessing and dataset splits achieved performance comparable to WaveNet on several classes, demonstrating that a non-autoregressive convolutional architecture can capture iEEG temporal patterns effectively. The specialized CNN reported by Nejedly et al. [9] highlights the trade off of highly tuned models: they maximize performance on particular signal types but require patient-specific tuning. WaveNet provides strong, consistent performance across classes, including improved pathological classification relative to the generalized CNN (F1 0.81 vs. 0.64). Which might reflect its deeper temporal receptive field and residual connections that capture long-range or irregular dynamics in iEEG signals.